OpenStudy (anonymous):

The function f(t) = 4t^2 - 8t + 7 shows the height from the ground f(t), in meters, of a roller coaster car at different times t. Write f(t) in the vertex form a(x - h)^2 + k, where a, h, and k are integers and interpret the vertex of f(t).

OpenStudy (anonymous):

f(t) = 4(t - 1)^2 + 3; the minimum height of the roller coaster is 3 meters from the ground f(t) = 4(t - 1)^2 + 3; the minimum height of the roller coaster is 1 meter from the ground f(t) = 4(t - 1)^2 + 2; the minimum height of the roller coaster is 2 meters from the ground f(t) = 4(t - 1)^2 + 2; the minimum height of the roller coaster is 1 meter from the ground